OpenStudy (anonymous):

How to arrive at the following from given \( x = r\sin \theta \cos \phi, y = r\sin \theta \sin \phi, z=r\cos\theta \) \[ \begin{bmatrix} A_x\\ A_y\\ A_z \end{bmatrix} = \begin{bmatrix} \sin \theta \cos \phi & \cos \theta \cos \phi & -\sin\phi\\ \sin \theta \sin \phi & \cos \theta \sin \phi & \cos\phi\\ \cos\theta & -\sin\theta & 0 \end{bmatrix} \begin{bmatrix} A_r\\ A_\theta\\ A_\phi \end{bmatrix} \] Also how show that \[ \begin{bmatrix} \hat i\\ \hat j\\ \hat k \end{bmatrix} = \begin{bmatrix} \sin \theta \cos \phi & \cos \theta \cos \phi & -\sin\phi\\ \sin \theta \sin \phi & \cos \theta \sin \phi & \cos\phi\\ \cos\theta & -\sin\theta & 0 \end{bmatrix} \begin{bmatrix} \hat e_r\\ \hat e_\theta\\ \hat e_\phi \end{bmatrix} \] How to change \( (a,b,c) \) into spherical polar coordinates and \( (r ,\theta, \phi) \) into Cartesian coordinates using this matrix?

OpenStudy (turingtest):

did you not drop an r from some parts of your matrix?

OpenStudy (unklerhaukus):

\[x=r\sin\theta\cos\phi\qquad\qquad y=r\sin\theta\sin\phi\qquad\qquad z=r\cos\theta\] \[x_r=\sin\theta\cos\phi\qquad x_\theta=r\cos\theta\cos\phi\qquad x_\phi=-r\sin\theta\sin\phi\]\[y_r=\sin\theta\sin\phi\qquad y_\theta=r\cos\theta\sin\phi\qquad y_\phi=r\sin\theta\cos\phi\]\[z_r=\cos\theta\qquad\qquad z_\theta=-r\sin \theta\qquad\qquad z_\phi=0\] \[\frac{\partial(x,y,z)}{\partial ( r,\theta, \phi)}=\left|\begin{matrix} x_r&x_\theta&x_\phi\\{y}_{r}&y_{\theta }&y_{\phi}\\{z}_{r}&{ z}_{\theta }&{z}_{\phi}\end{matrix}\right|\]

OpenStudy (unklerhaukus):

does this answer you question @hash.nuke do you need help calculating the Jacobian determinant?

OpenStudy (anonymous):

http://www.web-formulas.com/Math_Formulas/Linear_Algebra_Transform_from_Cartesian_to_Spherical_Coordinate.aspx @UnkleRhaukus I don't this is Determinant of Jacobian

OpenStudy (unklerhaukus):

pardon?

OpenStudy (turingtest):

I don't know what that site is @hash.nuke , but @UnkleRhaukus has given the correct Jacobian http://en.wikipedia.org/wiki/Spherical_coordinate_system#Cartesian_coordinates

OpenStudy (anonymous):

No ... that was just transformation rule for unit vectors. As there is no linear relation between curvilinear coordinate system and rectangular coordinate system, the transformation cannot be represented by matrix. I have come to conclude that.

OpenStudy (unklerhaukus):

\[\frac{\partial(x,y,z)}{\partial ( r,\theta, \phi)}=\left|\begin{matrix} x_r&x_\theta&x_\phi\\{y}_{r}&y_{\theta }&y_{\phi}\\{z}_{r}&{ z}_{\theta }&{z}_{\phi}\end{matrix}\right|\] \[=x_r\left((y_\theta\times z_\phi)-(y_\phi\times z_\theta)\right)-x_\theta\left((y_r\times z_\phi)-(y_\phi \times z_r)\right)+x_\phi\left((y_r\times z_\theta)-(y_\theta \times z_r)\right)\]\[=x_r\left((r\cos\theta\sin\phi\times 0)-(r\sin\theta\cos\phi\times -r\sin\theta )\right)\]\[ \qquad\qquad-x_\theta\left((\sin\theta\sin\phi\times 0)-(r\sin\theta\cos\phi \times \cos\theta)\right)\]\[\qquad\qquad+x_\phi\left((\sin\theta\sin\phi\times -r\sin\theta )-(r\cos\theta\sin\phi \times \cos\theta)\right)\] \[=x_r\left(0+r^2\sin^2\theta\cos\phi)\right)-x_\theta\left(0-r\sin\theta\cos\phi\cos\theta\right)-x_\phi\left(r\sin^2\theta\sin\phi+r\cos^2\theta\sin\phi\right)\]\[=x_rr^2\sin^2\theta\cos\phi+x_\theta r\sin\theta\cos\phi\cos\theta-x_\phi r\sin\phi\]\[=\sin\theta\cos\phi\ r^2\sin^2\theta\cos\phi+r\cos\theta\cos\phi\ r\sin\theta\cos\phi\cos\theta+r\sin\theta\sin\phi r\sin\phi\]\[=r^2\sin\theta\left(\sin^2\theta\cos^2\phi-\cos^2\theta\cos^2\phi+\sin^2\phi\right)\]\[=r^2\sin\theta\left((\sin^2\theta+\cos^2\theta)\cos^2\phi+\sin^2\phi\right)\]\[=r^2\sin\theta\left(\sin^2\phi+\cos^2\phi\right)\]\[{=r^2\sin\theta}\]\[\text{---------}\]